Write an equation in slope-intercept form for the line that passes through the given point and is parallel to the graph of the given equation
(3,-2), y = x + 4

Answer by TimothyLamb(4379)

parallel lines have equal slopes:
slope of new line = 1
---
y + 2 = 1(x - 3)
y + 2 = x - 3
y = x - 3 - 2
y = x - 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php